# Quantum Knowledge Distillation for Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) are integral to advancing natural language processing, used extensively from machine translation to content creation. However, as these models scale to billions of parameters, their resource demands increase dramatically. Meanwhile, quantum computing is recognized for efficiently solving complex problems with quantum characteristics like superposition and entanglement, providing a novel approach to these challenges. This paper attempts to combine quantum computing with LLMs and proposes a Quantum knowledge Distillation algorithm for LLMs (QD-LLM), aimed at reducing the computational and memory overhead required for model loading and inference. Specifically, during the distillation stage, data is fed simultaneously into both the LLMs and the designed quantum student model to initially quantify the difference between their outputs; subsequently, with the help of the true label, the optimization of the quantum student model is executed to minimize the difference with the LLM’s output. Throughout this process, only the parameters of the quantum student network are updated to make its output closer to that of the LLMs, thereby achieving the purpose of distillation. Finally, the optimized student model obtained by QD-LLM can efficiently solve domain-specific tasks during inference without the usage of the original LLMs. Experimental results show that, compared to mainstream compression methods, QD-LLM significantly reduces the number of training parameters, memory consumption, training time, and inference time while maintaining performance. Moreover, the optimized student model obtained by QD-LLM surpasses specific models designed for these tasks. We believe that QD-LLM can lay the groundwork for exploring the utilization of quantum computing in model compression and its potential extension to other natural language processing challenges.
keywords:
LLMs , Knowledge Distillation , Compression , Quantum Neural Networks , Quantum Computing
1 Introduction
The field of artificial intelligence has seen rapid advancements over the past years, leading to the emergence of numerous Large Language Models (LLMs) [1, 2, 3, 4]. These models, fueled by vast amounts of big data [5, 6, 7], have demonstrated significant advantages in language understanding across a wide range of domains, achieving consistently superior performance.
Utilizing open-source LLMs [8, 9, 10, 11] for various tasks requires loading the models for inference. However, the process of loading and running inference on LLMs incurs substantial computational and memory overhead, posing significant challenges for low-resource environments. To address this issue, extensive research has been conducted on model compression, aiming to significantly reduce the inference cost while minimizing the impact on model performance. These works can be broadly categorized based on their focus and underlying mechanisms into knowledge distillation [12, 13, 14], quantization [15, 16], and pruning [17, 18]. Knowledge distillation effectively compresses LLMs for specific tasks by designing student models and strategies [19, 20], or by fine-tuning student models directly using the outputs of teacher models [21], allowing student models to learn from the larger-scale networks. Quantization focuses on reducing the precision of model weights or activations through techniques such as matrix decomposition [15], second-order approximation [22], dynamic weight pruning [23], and adaptive precision allocation [24], effectively lowering the computational resources required and speeding up model inference [16]. Pruning reduces the complexity of large models by identifying and removing unimportant parameters [17] or modules [25], thus making the model more sparse and computationally efficient [26]. These methods have significantly reduced the computational burden of open-source LLMs while maintaining inference capability.
As LLMs continue to grow in size, there remains room for further optimization in model compression, especially in scenarios with extremely low resources or where faster inference is required. Traditional methods may struggle to meet these demands. Fortunately, extensive research has demonstrated that quantum computing can solve certain challenging problems efficiently using fewer quantum parameters [27, 28, 29, 30, 31]. This has led to the rise of Quantum Neural Networks (QNNs) [32, 33, 34], which show great potential in tasks such as classification [35] by leveraging quantum circuits to make efficient predictions with fewer parameters.
In this paper, we innovatively apply quantum computing to the distillation of LLMs and propose a novel Quantum knowledge Distillation algorithm for LLMs (QD-LLM), which utilizes the QNN to compress LLMs. In this method, a QNN designed specifically for large model knowledge distillation is proposed as the student model to distills knowledge from an LLM teacher model. During the training phase, the teacher model guides the training of the quantum circuit, allowing the predictions of the quantum circuit to gradually approach those of the teacher model. During the entire distillation process, the parameters of the quantum student network will be updated to align its output more closer with that of the LLMs, while the parameters of the LLMs remain unchanged, thus facilitating effective knowledge distillation. In the inference phase, the optimized student model obtained by QD-LLM operates independently of the LLMs, utilizing only the trained quantum circuit for efficient inference. Experiments on various tasks, such as multi-class topic analysis and binary sentiment analysis, show that QD-LLM significantly compresses LLMs while maintaining strong performance, resulting in faster inference speed and reduced memory consumption. This could potentially open new avenues for more efficient inference in language models. It should be noted that our experiments were conducted only on multiple classification tasks, as current computational capabilities do not support large-scale qubit simulations. So it is difficult to show similar advantages on complex tasks such as generation like LLMs.
The main contributions are summarized as follows.
  • 1.
    We propose a quantum computing-based model compression algorithm, termed the QD-LLM method. Distinct from traditional distillation methods that primarily rely on neural networks, this method applies quantum computing to the distillation of LLMs. The advantage of quantum computing lies in its ability to retain more information through the superposition and entanglement of quantum states. This enables the QD-LLM method to accelerate the distillation process while effectively enhancing the model’s capacity to process and preserve information.
  • 2.
    We designed a loss function suitable for QD-LLM. This loss function integrates the gap between the quantum student network with the LLM teacher network and the true label, providing a direct basis for the quantum student network to better learn effective knowledge.
  • 3.
    We conducted extensive experiments, and the results show that using QD-LLM to distill LLMs is feasible and demonstrates superior performance. Compared with knowledge distillation baselines, QD-LLM requires significantly fewer parameters, which directly leads to significantly reduced memory usage, training, and inference time. At the same time, the optimized quantum student network inherits the advantages of the LLMs and outperforms related-task baselines in terms of performance and inference time.
2 Related Work
2.1 Large language models
LLMs [4, 36] are equipped with a vast number of parameters that enable them to comprehend text from extensive datasets and tackle various Natural Language Processing (NLP) tasks. Recent advancements in LLMs have been made through techniques like instruction fine-tuning [37, 38] and reinforcement learning via human feedback [39, 40], significantly enhancing their performance. At the same time, many researchers are focusing on building open-source LLMs [8, 9, 10, 41], to drive advancements in specific domains and tasks. LLMs are primarily composed of encoder layers and multiple decoders based on the Transformer architecture [42]. Each decoder consists of an attention layer (Attn) and a feed-forward network (FFN). Given an input $T$, the model first processes it through the encoder to generate the corresponding vector representation $h_{0}$, followed by several decoders. The abstraction of this process is summarized as follows:
${h_{l}}={\text{FFN}}({h_{n^{\prime}}})+{h_{n^{\prime}}}={\text{FFN}}({\text{Attn}}({h_{n-1}})+{h_{n-1}})+{\text{Attn}}({h_{n-1}})+{h_{n-1}},$ (1)
where, $n$ represents the $n-th$ layer decoder,$n=1,2,\cdots,N$ and $N$ is the number of decoders contained in LLMs.
2.2 Quantum neural networks
QNNs, hybrid models that integrates parameterized quantum circuits with classical optimizers, leverages the unique capabilities of quantum mechanics to enhance computational performance and optimization efficiency. These networks have gained momentum in various machine learning tasks, particularly in classification, due to their ability to exploit quantum characteristics for handling complex datasets. Bausch [43] et al. introduced the first quantum recurrent neural network, demonstrating its effectiveness in sequence learning and classification tasks. Du [44] et al. systematically investigated the problem-dependent power of quantum neural classifiers (QCs) on multi-class classification tasks, revealing that QCs exhibit a U-shaped risk curve and can outperform classical classifiers in certain tasks, such as parity datasets. In text classification, where a significant work is BERT Quantum Text Classifier (BERT-QTC) [45]. BERT-QTC demonstrates how quantum classifiers can enhance performance in high-dimensional feature spaces, achieving superior accuracy compared to classical classifiers in NLP tasks. Our work also draws inspiration from this model, especially how it combines quantum circuits with classical neural networks to improve performance. We referenced some of their designs for quantum layers and variational circuits when designing the quantum part of our own hybrid model.
2.3 Knowledge distillation
Despite the remarkable performance of LLMs, they often come with substantial computational and storage demands. As a result, model compression is essential for the practical deployment of these models on low-resource machines and for further research. One of the mainstream approaches for model compression is Knowledge Distillation [12], which compresses a large teacher model by training a smaller student model under the teacher’s guidance [13]. In computational linguistics, this method either uses the teacher model’s outputs at each time step as guidance [19] or fine-tunes the student model directly using the outputs from the teacher model [21, 41]. At present, some quantum distillation algorithms have been proposed. For example, Hasan et al. [46] and Li et al. [47] used the classical convolutional neural network (CNN) and classical machine learning as teachers to guide the training of the quantum student model, thereby significantly improving the performance of the quantum model; Alam et al. [48] proposed a method for enhancing QNNs through knowledge distillation. However, these algorithms are mainly targeted at small-scale models such as CNN, and have not yet explored the compression application of LLMs. In this era of explosive development of large models, the issue of using quantum computing to compress LLMs is worth exploring.
3 Preliminary
3.1 Quantum computing: fundamental concepts
Quantum computing is based on the principles of quantum mechanics and offers significant advantages over classical computing in terms of parallel processing and solving complex problems. Classical computers use bits as the smallest unit of information, where each bit can only exist in one of two states: 0 or 1. In contrast, quantum computers use quantum bits (qubits), which can exist in a superposition of states, allowing them to process multiple solutions simultaneously. The state of a quantum bit can be expressed as a quantum superposition:
$|\psi\rangle=\alpha|0\rangle+\beta|1\rangle,$ (2)
where, $\alpha$ and $\beta$ are complex coefficients that satisfy the normalization condition:
$|\alpha|^{2}+|\beta|^{2}=1.$ (3)
This equation implies that the quantum bit exists in a superposition of both $|0\rangle$ and $|1\rangle$ states, with $|\alpha|^{2}$ and $|\beta|^{2}$ representing the probability amplitudes of measuring the quantum bit in either of the states.
The power of quantum computing lies in its ability to utilize quantum entanglement and quantum interference. Quantum entanglement refers to the phenomenon where multiple qubits are correlated such that the state of one qubit cannot be described independently of the state of the other qubits, even when they are physically separated. Quantum algorithms such as Shor algorithm and Grover algorithm exploit the parallelism provided by quantum superposition and entanglement, enabling them to outperform classical algorithms in specific problem domains. Despite the enormous potential, quantum computing still faces many challenges, particularly related to qubit scalability, noise management, and error correction. Nevertheless, it is widely regarded as having the potential to revolutionize fields such as optimization, material science, drug discovery, and beyond. This capability enables quantum computers to evaluate many possible solutions in parallel, offering exponential speedup over classical counterparts in certain computational tasks. As a result, quantum computing holds the potential to solve complex problems that are infeasible for classical computers, particularly in fields such as optimization, cryptography, and simulation.
3.2 Quantum gates
Quantum computation relies on quantum circuits, and the core operations within these circuits are quantum gates. Quantum gates not only modify the states of quantum bits but also influence the computation process via quantum interference effects. Similar to classical logic gates in conventional computing, quantum gates are the fundamental building blocks of quantum computing. By combining different quantum gates, it is possible to perform complex computational tasks. Quantum circuits typically consist of multiple quantum gates. These gates, through entanglement and superposition of qubits, perform computations and optimize model parameters.
3.2.1 Common quantum gates
Pauli gates are a set of fundamental quantum gates, including Pauli-X, Pauli-Y, and Pauli-Z gates. They perform rotation operations on quantum bits, altering their states. These gates are widely used in QNNs, especially in the control of quantum states and the generation of quantum entanglement.
Pauli-X Gate (X Gate)
The X gate is equivalent to the classical NOT gate and swaps the states $|0\rangle$ and $|1\rangle$. The matrix representation of the X gate is:
$X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}.$ (4)
In QNNs, the X gate is used to perform basic bit-flipping operations, often employed in initializing and transforming quantum states.
Pauli-Z Gate (Z Gate)
The Z gate applies a phase flip to the quantum bit. It leaves $|0\rangle$ unchanged, but it introduces a negative phase when applied to $|1\rangle$. The matrix representation is:
$Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$ (5)
The Z gate is used in quantum computing for phase rotations and the control of quantum entanglement, where it selectively flips the phase of a quantum state.
Pauli-Y Gate (Y Gate)
The Y gate performs a complex rotation, transforming $|0\rangle$ to $-i|1\rangle$ and $|1\rangle$ to $i|0\rangle$. The matrix representation is:
$Y=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}.$ (6)
3.2.2 Hadamard Gate (H Gate)
The Hadamard gate is a crucial quantum gate, especially in QNNs. It transforms quantum bits from definite states (e.g., $|0\rangle$ or $|1\rangle$) into uniform superposition states. Specifically, the Hadamard gate maps both $|0\rangle$ and $|1\rangle$ to superposition states, allowing quantum bits to be in multiple states simultaneously. This superposition effect enables parallel computation across multiple potential states, providing significant computational power for quantum circuits. The effect of the Hadamard gate can be described as follows:
$H|0\rangle=\frac{1}{\sqrt{2}}(|0\rangle+|1\rangle),\quad H|1\rangle=\frac{1}{\sqrt{2}}(|0\rangle-|1\rangle).$ (7)
Its matrix representation is:
$H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}.$ (8)
The Hadamard gate is used to place quantum bits into superposition, allowing quantum computation to process multiple potential states in parallel. In QNNs, the Hadamard gate is commonly used to generate complex quantum states, providing richer representations for model optimization and training, especially when dealing with high-dimensional data and solving complex optimization problems.
3.2.3 CNOT Gate (Controlled-NOT Gate)
The Controlled-NOT (CNOT) gate is a two-qubit gate, widely used in quantum computing and QNNs, particularly in the creation of quantum entanglement. The CNOT gate works by flipping the target qubit when the control qubit is in the state $|1\rangle$; otherwise, the target qubit remains unchanged. The matrix representation of the CNOT gate is:
$\text{CNOT}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}.$ (9)
In QNNs, the CNOT gate is primarily used to create entanglement between quantum bits, which is a unique feature of quantum computing. Quantum entanglement allows QNNs to represent more complex relationships between data and enables more efficient computations during model training.
3.2.4 CZ Gate (Controlled-Z Gate)
The Controlled-Z (CZ) gate is also a two-qubit gate. It applies a Z transformation to the target qubit when the control qubit is in the state $|1\rangle$, introducing a negative phase. Its matrix representation is:
$\text{CZ}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&-1&0\\ 0&0&0&1\end{pmatrix}.$ (10)
The CZ gate is commonly used in QNNs for phase control between quantum bits. It plays a crucial role in adjusting the quantum states’ phases and helps quantum networks capture data features across higher dimensions. In quantum machine learning and quantum optimization, the CZ gate is used to adjust quantum states, improving performance in high-dimensional spaces.
3.3 Simulation
Currently, quantum computers are in the Noisy intermediate-scale quantum (NISQ) era because they cannot effectively simulate or perform large-scale, high-fidelity quantum computing tasks due to their limited number of qubits, insufficient error correction capabilities, and high noise levels. In this era, the most existing quantum algorithms still rely on classical computers to simulate quantum circuits.
To simulate quantum algorithms in classical computers, researchers have developed Python packages including mindquantum, Qiskit, Cirq, and Torchquantum, which support frameworks that allow researchers to simulate of quantum gates and circuits in quantum networks, making them key players in model optimization and training.
4 Methodology
4.1 QD-LLM Overall
@@
4.2 Quantum distillation for LLMs
4.2.1 LLM teacher model
@@
The function $LM(\cdot)$ abstracts formula (1) and represents the complete language model calculation process, where ${\text{Softmax}}(\cdot)$ represents the probability conversion process. Constructing the vocabulary $V$ is central to the operation of LLMs, and the order and size of vocabularies differ between models. To distill a generative LLM effectively, the student model must align with the teacher model’s vocabulary. However, current vocabularies have reached the scale of tens of thousands, which cannot be effectively learned by the limited number of qubits in quantum models. Therefore, in this paper, we mainly study the distillation of LLMs on multiple classification tasks (e.g., sentiment analysis, subject analysis, etc.) In these tasks, we convert generative LLMs into a classification mode, i.e. given input data $x$, the LLMs output the probability $P(y|x)$ that $x$ belongs to a certain class, represented as a feature vector. This transformation removes the dependency on large vocabularies, allowing effective distillation into quantum student models.
For LLMs to perform optimally on specific domains and tasks, fine-tuning is necessary using domain-specific data. However, full parameter fine-tuning of LLMs would lead to prohibitive computational costs. Therefore, this work employs the Low-Rank Adaptation (LoRA) [49] technique for efficient fine-tuning, which means that LoRA here is just to get the output of LLM teachers. LoRA modifies only a subset of the LLM’s weights by adding a low-rank $\Delta{\mathbf{w}}$ matrix to the original weight matrix ${{\mathbf{W}}_{0}}\in{\mathbb{R}^{d\times k}}$, $\Delta{\mathbf{w}}$ is determined by two low-rank (i.e., low-dimensional) matrices $a$ and $b$, with the following relationship:
@@
where, $\circ$ represents the matrix outer product, $r$ is the rank of $a$ and $b$ , $d$ and $k$ are the linear layer dimensions and number of attention heads of the original LLMs, respectively. When resources are limited or the pre-trained knowledge is required to retain, LoRA provides a method to efficiently fine-tune large models. Finally, the original LLMs parameters and $\Delta{\mathbf{w}}$ are merged to obtain the final fine-tuned LLMs parameters ${{\mathbf{W}}_{ft-LLM}}$:
@@
In this way, the complete LLMs to be distilled are obtained. That is, LoRA only affects the output of LLM, and the fine-tuned LLM is what really needs distillation, as shown in Formula 13.
4.2.2 Quantum student model
The student network is composed of a classical-quantum data encoding layer and an ansatz with $p$ layers, which harness the advantages of quantum circuits in processing complex data. The classical-quantum encoding layer retains the semantic information from the classical vector, ensuring an accurate representation of the classical semantics. Meanwhile, the ansatz introduces trainable parameters, enhancing the expressiveness and flexibility of the quantum circuit. This design enables the quantum student network to capture complex semantic information from text, while leveraging quantum state evolution in high-dimensional space to improve the learning and classification capabilities of the student model.
Classical-Quantum Data Coding Layer
For the quantum circuit to process textual data, the text needs to be encoded into quantum states. Previous research [50] showed that one-hot encoding can map each token to a distinct initial quantum state, suitable for small vocabulary tasks. However, since each qubit can only represent a limited number of states, one-hot encoding becomes impractical for tasks involving large vocabularies.
In QD-LLM, an embedding layer [51] is used to map token sequences into high-dimensional semantic vectors E. Please note: the embedding layer is not trainable and serves solely as a mapping tool. The semantic vectors are then passed through a fully connected layer for dimensionality reduction, ensuring that the final vector dimension matches the number of qubits. The dimensionality reduction is expressed by the following formula:
@@ 4.2.2 Quantum student model
Figure 2: The structure of the parameterized quantum circuit. It presents the architecture of a complete variational quantum circuit, including both the encoding layer for classical-to-quantum data transformation and the Ansatz component. Figure 2 a) represents the encoding layer, where classical input data is encoded into quantum states through parameterized single-qubit rotation gates $R_{X}$. Figure 2 b) illustrates the Ansatz, which consists of multiple layers of parameterized rotation gates and entangling gates, used to further optimize the quantum states based on the training parameters. The overall circuit improves the expressiveness of the quantum model by stacking rotation and entanglement operations, enabling it to better capture the features of the input data.
Finally, the output of the quantum measurement is passed through a softmax function to produce the output of the student model.
@@
$|{\psi_{{\text{enc}}}}\rangle=\prod\limits_{i=1}^{n}{{R_{X}}}({\theta_{i}})|0{\rangle^{n}},$ (16)
The initial quantum state $|0{\rangle^{n}}$ represents all qubits in the $|0{\rangle}$. At this point, the classical vector has been successfully mapped into the quantum state $|{\psi_{{\text{enc}}}}\rangle$ preparing it for further quantum operations.
Ansatz Design
In the parameterized quantum circuit Ansatz designed by QD-LLM, we enhance the expressiveness of quantum circuits through a series of quantum gate operations, enabling the model to better capture and process complex input information. The parameters of these gates are optimized during the training process, thereby improving the circuit’s ability to represent the features of the input data. The structure of the Ansatz is illustrated in Figure 2 b).
For a single layer of the Ansatz, each qubit $q_{i}$ first undergoes a unified rotation operation $U_{Y}(\theta_{i})$, consolidating the effects of three consecutive $R_{Y}$ gates into a single more potent operation. This enhances the qubit states’ ability to flexibly express the input data features. The corresponding equation for $U_{Y}$ is given as follows:
${U_{Y}}\left({{\phi_{i}}}\right)={R_{Y}}\left({{\phi_{i1}}}\right)+{R_{Y}}\left({{\phi_{i2}}}\right)+{R_{Y}}\left({{\phi_{i3}}}\right)$ (17)
@@
where $\phi_{i}=\phi_{i1}+\phi_{i2}+\phi_{i3}$ is the cumulative rotation angle for qubit $q_{i}$. Here, $\phi_{i1},\phi_{i2},$ and $\phi_{i3}$ are the rotation angles corresponding to each of the three applications of the $R_{Y}$ gate, and $\sigma_{Y}$ is the Pauli-Y matrix.
@@
where $\delta_{i,j}$ is the entanglement parameter between qubits $q_{i}$ and $q_{j}$, and $\sigma_{Z}$ is the Pauli-Z matrix.
@@
where $\lambda_{i}=\lambda_{i1}+\lambda_{i2}+\lambda_{i3}$ is the cumulative rotation angle for qubit $q_{i}$. Here, $\lambda_{i1},\lambda_{i2},$ and $\lambda_{i3}$ are the rotation angles corresponding to each of the three applications of the $R_{Z}$ gate.
@@
This operation conditions the state of $q_{j}$ on $q_{i}$, applying an $X$ gate (flip operation) to $q_{j}$ if $q_{i}$ is in the $|1\rangle$ state, and no operation if $q_{i}$ is in the $|0\rangle$ state.

### v2

@@ Abstract
As foundational tools in natural language processing, Large Language Models (LLMs) have immense parameter scales, which makes deployment and inference increasingly prohibitive, especially in resource-constrained devices. Therefore, knowledge distillation for LLMs, i.e., compressing the LLM to a smaller model, is meaningful. With strong parameter representation capacity, quantum computing is regarded as a promising solution. Here, we propose a Quantum knowledge Distillation model for LLMs (QD-LLM) that leverages variational quantum circuits to learn from LLMs. In classical simulation, QD-LLM outperforms several mainstream distillation methods on multiple text classification tasks in terms of both accuracy and efficiency using only 11 qubits. The results reveal an interesting phenomenon that the simulation of quantum student models may be regarded as a new class of quantum-inspired classical algorithms. Remarkably, we deploy the obtained circuits on the Baihua superconducting quantum processor via the Quafu platform to assess practical feasibility. The model maintains stable inference performance despite hardware constraints such as decoherence and finite sampling. In summary, QD-LLM marks a foundational step in connecting quantum computing with LLMs, demonstrating the feasibility of quantum-native approaches that aim to compress and deploy models of increasingly larger scales. The code of this article has been open-sourced at https://github.com/Lilingxiao-bupt/QD-LLM.
Introduction
Recent advances in artificial intelligence have led to the emergence of numerous Large Language Models (LLMs) 1,2,3,4, which have become the backbone of modern natural language processing (NLP) systems, consistently delivering state-of-the-art performance across a wide range of tasks 5,6,7. However, deploying LLMs 8,9,10,11 for various tasks typically requires loading the models for inference, which results in substantial computational and memory overhead. To address this problem, Knowledge Distillation (KD) for LLMs 12,13 has been widely adopted as a means of transferring task-specific knowledge from large-scale models (called teacher models) to smaller models (called student models) 14. Distilled models offer lower computational costs compared to using LLMs directly while preserving comparable inference capability.
As LLMs continue to grow in parameter count, there remains room for further optimization in model compression, especially in scenarios with extremely low resources or where faster inference is required. Fortunately, extensive research has demonstrated that quantum computing can solve problems efficiently using fewer quantum parameters 15,16,17,18,19 due to its powerful parameter representation capability. In particular, quantum neural networks (QNNs)  20,21,22,23 have been explored as student models in knowledge distillation, with most existing studies focusing on small-scale image classification tasks using shallow classical 24,25 or quantum teacher models 26. However, current QNN-based student models have not yet demonstrated the ability to distill LLMs for natural language classification tasks. In other words, whether these models can capture the complexity, semantic richness, and real-world variability inherent in natural language understanding of LLMs remains an open question.
To bridge this challenge, we propose a Quantum knowledge Distillation algorithm for LLMs (QD-LLM). Unlike previous quantum distillation studies limited to small models or small tasks, QD-LLM fills the gap of LLMs distillation by treating LLMs as fixed teacher models and training a quantum student model to approximate their outputs. The student model is implemented as a variational quantum circuit specifically designed for this distillation setting. During the training phase, the circuit parameters are optimized to minimize the difference between the outputs of the student model and the teacher model. Once trained, the student performs inference independently, relying solely on the learned circuit. This makes it well-suited for deployment in scenarios with limited computational resources. We evaluate QD-LLM through a comprehensive series of simulation experiments and real-device deployments, where the model is implemented as a variational quantum circuit with only 11 qubits and assessed on three real-world natural language processing tasks: Emotion Analysis 27, Hiding Detection 28, and Thematic Analysis 29. To ensure a fair comparison, all models, including quantum models simulated on classical hardware and classical baselines, are trained under identical conditions with aligned input representations. Compared to mainstream classical baselines, QD-LLM significantly reduces memory consumption and achieves efficient training and inference on classification tasks due to lower parameter counts. In addition, we deploy the quantum model obtained from QD-LLM in a real quantum device, the Baihua processor provided by the Quafu platform 30, to examine the practical feasibility of deploying QD-LLM. Our student model achieves stable and reliable inference results in real devices, despite hardware constraints such as decoherence and shot noise.
Interestingly, our simulation results reveal a compelling phenomenon: the quantum student model obtained from our approach outperforms existing mainstream classical models, including the ones distilled from LLM and the ones from task-specific training, in both memory consumption and classification accuracy on practically meaningful tasks. This unexpected observation reveals that the simulation of our quantum student model can be regarded as a new class of quantum-inspired classical algorithm, which is a direct simulation of the quantum algorithm. Furthermore, the strong simulation results imply that, as the quality of qubits in real hardware continues to improve, our quantum student model may emerge as a promising candidate for achieving quantum supremacy on practically meaningful tasks, even though the inference accuracy on current hardware is limited. Note that the above observations benefit from the fact that only a small-scale quantum circuit (here, 11 qubits and a shallow depth) is enough for the above tasks in our student method.
Results
QD-LLM Overall
@@ Results
Quantum distillation for LLMs
LLM teacher model
@@
[…]
For LLMs to perform optimally on specific domains and tasks, fine-tuning is necessary using domain-specific data. However, full parameter fine-tuning of LLMs would lead to prohibitive computational costs. Therefore, this work employs the Low-Rank Adaptation (LoRA) 31 technique for efficient fine-tuning, which means that LoRA here is just to get the output of LLM teachers. LoRA modifies only a subset of the LLM’s weights by adding a low-rank $\Delta{\mathbf{w}}$ matrix to the original weight matrix ${{\mathbf{W}}_{0}}\in{\mathbb{R}^{d\times k}}$, $\Delta{\mathbf{w}}$ is determined by two low-rank (i.e., low-dimensional) matrices $a$ and $b$, with the following relationship:
@@
where, $\circ$ represents the matrix outer product, $r$ is the rank of $a$ and $b$ , $d$ and $k$ are the linear layer dimensions and number of attention heads of the original LLMs, respectively. When resources are limited or the pre-trained knowledge is required to be retained, LoRA provides a method to efficiently fine-tune large models. Finally, the original LLMs parameters and $\Delta{\mathbf{w}}$ are merged to obtain the final fine-tuned LLMs parameters ${{\mathbf{W}}_{ft-LLM}}$:
@@
In this way, the complete LLMs to be distilled are obtained. That is, LoRA only affects the output of LLM, and the fine-tuned LLM is what really needs distillation, as shown in Formula 3.
Quantum student model
The student network is composed of a classical-quantum data encoding layer and an ansatz with $p$ layers, which harnesses the advantages of quantum circuits in processing complex data. The classical-quantum encoding layer retains the semantic information from the classical vector, ensuring an accurate representation of the classical semantics. Meanwhile, the ansatz introduces trainable parameters, enhancing the expressiveness and flexibility of the quantum circuit. This design enables the quantum student network to capture complex semantic information from text while leveraging quantum state evolution in high-dimensional space to improve the learning and classification capabilities of the student model. While this circuit serves as an effective demonstration of the framework, it is not necessarily the optimal architecture. The QD-LLM framework accommodates alternative ansatz designs, and future exploration of deeper or hardware-specific quantum circuits may yield further improvements.
Classical-Quantum Data Coding Layer. For the quantum circuit to process textual data, the text needs to be encoded into quantum states. Previous research 32 showed that one-hot encoding can map each token to a distinct initial quantum state, suitable for small vocabulary tasks. However, since each qubit can only represent a limited number of states, one-hot encoding becomes impractical for tasks involving large vocabularies.
In QD-LLM, an embedding layer 33 is used to map token sequences into high-dimensional semantic vectors E. Please note: the embedding layer is not trainable and serves solely as a mapping tool. The semantic vectors are then passed through a fully connected layer for dimensionality reduction, ensuring that the final vector dimension matches the number of qubits. The dimensionality reduction is expressed by the following formula:
@@
Figure 2: The structure of the parameterized quantum circuit. It presents the architecture of a complete variational quantum circuit, including both the encoding layer for classical-to-quantum data transformation and the Ansatz component. Figure 2 a) represents the encoding layer, where classical input data is encoded into quantum states through parameterized single-qubit rotation gates $R_{X}$. Figure 2 b) illustrates the Ansatz, which consists of multiple layers of parameterized rotation gates and entangling gates, used to further optimize the quantum states based on the training parameters. The overall circuit improves the expressiveness of the quantum model by stacking rotation and entanglement operations, enabling it to better capture the features of the input data. Finally, the output of the quantum measurement is passed through a softmax function to produce the output of the student model.
@@
$|{\psi_{{\text{enc}}}}\rangle=\prod\limits_{i=1}^{n}{{R_{X}}}({\theta_{i}})|0{\rangle^{n}}.$ (6)
[…]
Ansatz Design. In the parameterized quantum circuit Ansatz designed by QD-LLM, we enhance the expressiveness of quantum circuits through a series of quantum gate operations, enabling the model to better capture and process complex input information. The parameters of these gates are optimized during the training process, thereby improving the circuit’s ability to represent the features of the input data. The structure of the Ansatz is illustrated in Figure 2 b). For a single layer of the Ansatz, each qubit $q_{i}$ first undergoes a unified rotation operation $U_{Y}(\theta_{i})$, consolidating the effects of three consecutive $R_{Y}$ gates into a single more potent operation. This enhances the qubit states’ ability to flexibly express the input data features. The corresponding equation for $U_{Y}$ is given as follows:
${U_{Y}}\left({{\phi_{i}}}\right)={R_{Y}}\left({{\phi_{i1}}}\right)+{R_{Y}}\left({{\phi_{i2}}}\right)+{R_{Y}}\left({{\phi_{i3}}}\right),$ (7)
@@
where, $\phi_{i}=\phi_{i1}+\phi_{i2}+\phi_{i3}$ is the cumulative rotation angle for qubit $q_{i}$. Here, $\phi_{i1},\phi_{i2},$ and $\phi_{i3}$ are the rotation angles corresponding to each of the three applications of the $R_{Y}$ gate, and $\sigma_{Y}$ is the Pauli-Y matrix.
@@
where, $\delta_{i,j}$ is the entanglement parameter between qubits $q_{i}$ and $q_{j}$, and $\sigma_{Z}$ is the Pauli-Z matrix.
@@
where, $\lambda_{i}=\lambda_{i1}+\lambda_{i2}+\lambda_{i3}$ is the cumulative rotation angle for qubit $q_{i}$. Here, $\lambda_{i1},\lambda_{i2},$ and $\lambda_{i3}$ are the rotation angles corresponding to each of the three applications of the $R_{Z}$ gate.
@@
This operation conditions the state of $q_{j}$ on $q_{i}$, applying an $X$ gate (flip operation) to $q_{j}$ if $q_{i}$ is in the |1|1$\rangle$| 1 ⟩ state, and no operation if $q_{i}$ is in the |0|0$\rangle$| 0 ⟩ state.